The weekly wages of employees of Volta gold are normally distributed about a mean of$1,250 with a standard deviation of $120. Find the probability of an employee having a weekly wages lying between$1,320 and $970


1
Expert's answer
2021-07-19T16:14:50-0400

"\mu=1250 \\
\sigma=120 \\
P(970<X<1320) = P(X<1320) -P(X<970) \\
= P(Z< \frac{1320-1250}{120}) -P(Z< \frac{970-1250}{120}) \\
= P(Z< 0.583) -P(Z< -2.333) \\
= 0.7200 -0.0098 \\
= 0.7102"